[Brahms-l] Shift report 20060620 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 20 2006 - 00:01:47 EDT
Shifter leader was SB
Shifters were 


<pre>
1. Beam from day shift continued. 
   Data taken throughout the evening shift.
   Beam scheduled to be dumped at 00:40
2. Runs 16261 - 16273, 16274 going on
   Total trig2 events = 278486
3. No new detector problems
</pre>
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 20 00:03:22 2006

This archive was generated by hypermail 2.1.8 : Tue Jun 20 2006 - 00:03:36 EDT